what reason would be used if the statement was m<1+m<2=m<3+m<4?
Answer by nyc_function(2733)

When the lines are parallel:

Interior Angles on the Same Side of the Transversal

(measures are supplementary)

Their "name" is simply a description of where the angles are located.

Interiors on the same side of the transversal make the following possible:

m<1 + m<2 = 180 degrees.

m<3 + m<4 = 180 degrees.

The conclusion is m<1 + m<2 = m<3 + m<4.